Haas Company manufactures and sells one product. The following information pertains to each of the companys first three years of operations: Variable costs per unit Manufacturings Direct material8 27 14 Direct labor Variable manufacturing overhead Variable selling and administrative Fixed costs per year Pixed manufacturing overhead 510,000 Pixed selling and administrative expenses 210,000 During its first year of operations, Haas produced 60,000 units and sold 60,000 units. During ts second year of operations, it produced 75,000 units and sold 50,000 units. In its third y price of the companys product is $60 per unit ear, Haas produced 40,000 units and sold 65,000 units. The selling Required 1. Compute the companys break-even point in unit sales. 2. Assume the company uses variable costing: a. Compute the unit product cost for Year 1, Year 2, and Year 3 b. Prepare an income statement for Year 1, Year 2, and Year 3 3. Assume the company uses absorption costing: a. Compute the unit product cost for Year 1, Year 2, and Year 3. b. Prepare an income statement for Year 1, Year 2, and Year 3.
Complete this question by entering your answers in the tabs below. Req 1 Compute the companys break-even point in unit sales. Break-even unit sales Req 2A Req 2B Req 3A Req 38 units Req 2A >
Complete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req 3A Req 38 Compute the unit product cost for Year 1, Year 2, and Year 3. Assume the company uses variable costing. Year 1 Year 2 Year 3 Unit product cost C Req 1 Req 2B >
Complete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req 3A Req 38 Prepare an income statement for Year 1, Year 2, and Year 3. Assume the company uses variable costing. Haas Company Variable Costing Income Statement Year 1 Year 2 Year 3 0 Net operating income (loss) Req 2A Req 3A
Complete this question by entering your answers in the tabs below. Req 1 Req 2A Req 28 Req 3A Req 3B Compute the unit product cost for Year 1, Year 2, and Year 3. Assume the company uses absorption costing. (Round your intermediate calculations and final answers to 2 decimal places.) Year 1 Year 2 Year 3 Unit product cos C Req 28 Req 3B
stuteeHt Tor Year 1, Year 2, and Year 3. Complete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req 3A Req 38 Prepare an income statement for Year 1, Year 2, and Year 3. Assume the company uses absorption costing. (Round your intermediate calculations to 2 decimal places.) Haas Company Absorption Costing Income Statement Year 1 Year 2 Year 3 Net operating income (loss) C Req 3A
Answer #1

Ans. 1 Calculation of Break even point

Selling price of the product           :      60

Direct material                        : 27                     

Direct labour                           : 14

Variable manufactured o/h      : 6

Variable selling and admn exp.: 1        : 48

Contribution p.u(60-48)                                 12

Calculation of total fixed cost

Fixed manufactured o/h   =    510000

Fixed Selling and admn exp.      =    210000

Total fixed cost                           =    720000

Break even point   = total fixed cost/contribution per unit

                              = 720000/12 = 60000 units

Ans. 2 if company using variable costing

a. calculation of unit cost under variable costing

         Direct material             :           27

         Direct labour                :          14

        Variable manufactured :        6

        Selling and admn cost :          1  

Total unit cost under variable cost $48, Variable cost will be same all the three years

                                                  INCOME STATEMENT UNDER VARIABLE COSTING

                                                                1 Yr                                   2Yr                          3yr

Sales     (a)                                         3600000                            3000000                  3900000

                                                       (60000X60)                        (50000X60)               (65000X60)

Less:Unit cost under variable cost 2880000    3600000                   1920000

(unit produced X48)              

Add: Opening Stock (unitX48)               -                                         -                            1200000*

Less: Closing stock (unit X48) -    1200000*    -

Cost of goods sold(b)                       2880000                              2400000                   3120000

Contribution margin(a-b)     720000     600000     780000

Less: Fixed overhead 510000     510000 510000

   Fixed selling and admn exp 210000    210000                     210000

Operating profit                                  0                                         (120000)                   60000

Note : calculation of closing stock for 2nd yr and opening stock of 3rd yr

unit produced in the year of 2 = 75000

Sold                                         = 40000

closing stock                            = 25000

this stock will be opening stock next yr

Ans 3. if company following Absorption costing

1. Calculation of unit cost under absorption costing

    Variable manufacturing overhead   (27+14+6) = $47

    Fixed manufacturing cost for 1yr = 510000/60000 = 8.5

   Fixed manufacturing cost for 2nd yr = 510000/75000 = 6.8

   Fixed manufacturing cost for 3rd yr = 510000/40000 = 12.75

Cost of unit for the year of 1 = 47+8.5 = $55.5

Cost of unit for the year of 2 = 47+6.8 = 53.80

Cost of unit for the year of 3 = 47+12.75 = 59.75

                                                              INCOME STATEMENT UNDER ABSORPTION COSTING

                                                                    YR 1                         YR 2                             YR3

Sales (same as variable) (a)                    3600000                     3000000                       3900000

Less: cost of production                      3330000                     4035000                       2390000

                                                            (55.5X60000)    (75000X53.80) (40000X59.75)   

Add: Opening stock                - -    1345000

                                                                                                                                      (25000X53.80)

Less: Closing stock                                    -                              (1345000)                            -

                                                                                                (25000X53.80)

Cost of goods sold (b)                            3330000                     2690000                          3735000

Gross profit (a-b) 270000 310000    165000

Variable selling exp(sold unitX1)     (60000)    (50000)    (65000)

Fixed selling exp (210000)    (210000)    (210000)

Operating profit/loss 0    50000                        (110000)